# High-Fidelity Diffusion-based Image Editing

### 25 Dec 2023 v1

@@ Introduction
Figure 2: Overview of our proposed rectifier framework. The rectifier is a hypernetwork consisting of a global encoder and multiple subnet branches. It takes as input the original image $\bm{x}_{0}$ and the estimation at each step ($\mathbb{P}_{t}[\bm{\epsilon}_{t}^{\theta}(\bm{x}_{t})]$), targets to modulate the degraded residual features into offset weights, providing compensated information for high-fidelity reconstruction. We select the middle and up-sampling blocks of U-Net for modulate, considering that these blocks contain both low-level semantic information and high-resolution details. We also employ separable convolution to reduce the amount of generated parameters.
@@
While these methods point out how to compensate for the information gap, their reconstructions are still far from high-fidelity, and resorting to external network or classifier also makes it difficult to generalize to semantic editing tasks which mainly relies on diffusion’s internal representations. In this work, we propose a framework where a rectifier is incorporated to modulate residual features into offset weights, providing compensated information to help pretrained diffusion model achieving high-fidelity reconstructions. The framework is illustrated in Fig. 2. Our rectifier is a hypernetwork (David, Andrew, and Quoc 2016) which takes as input every step’s estimation and original image, expected to exploit the degradated residual features to fill the fidelity gap. The inputs first pass through a global encoder, then transformed by a series of sub-nets to generate layer-wise modulation. We choose to modulate the middle and up-sampling blocks of U-Net (Ronneberger, Fischer, and Brox 2015), considering they contain both high-resolution and low-level semantic information. Furthermore, without the interference of other representations like classifier, our framework is highly adaptive for semantic editing tasks, and is easier to generalize to other diffusion-based downstream tasks.
@@
Another advantage of our editing training strategy is that we do not need to specify any heuristic-defined parameters to fit different attributes. It should be remarked here that for methods like Asyrp (Kwon, Jeong, and Uh 2022) and DiffusionCLIP (Kim, Kwon, and Ye 2022), neither of them employ editing through the entire denoising process. Asyrp halts editing prematurely and adds stochastic noise then to boost preceived quality, while DiffusionCLIP does not inverse images into complete noise for preserving their morphologies. Setting these parameters meticulously for every separate attribute is intricate and bothersome. Our method though, starts from pure noise and traverses the process throughly to get editing results, no extra settings are needed.
@@
In a comparable manner to reconstruction part, both quantitative and qualitative evaluation of editing are exhibited here too. Fig.5 presents some qualitative comparisons towards different methods trained under 50 inversion and sampling steps. Like previously noted, neither of Asyrp nor DiffusionCLIP employ edit through the entire denoising process, Asyrp applies stochastic noise during final process to boost perceived quality, and DiffusionCLIP begins editing from intermediate noisy images for preserving their morphologies. Our method, though starts from complete noise and traverses the whole denoising process to edit, still carries out editing results with exceptional quality. Several illustrative examples are, for instance, elements like hat and ring are kept intact during semantic editing, along with the preservation of image’s overall shape and background. Meanwhile, distortions or artifacts brought by conditional input text are avoided, and vital information loss is also alleviated.
@@
Doubts may arise regarding to whether our gain in editing comes from the rectifier or the training strategy. With the benefit of rectifier already proven in preceding part, we now focus on ablation studies to validate the effectiveness of our editing training strategy alone. Attribute ”smiling” is chosen for evaluating various algorithms. Fig.6 presents the performance of these algorithms. As shown, even without any enhancement from rectifier, our strategy still produces results with less distortion and information loss compared to others. This demonstrates our strategy could be employed as an independent and generalized training approach for editing tasks based on diffusion models. Furthermore, recall that neither Asyrp nor DiffusionCLIP implement editing through the entire process, they either stop prematurely or start from incomplete noise. We thus investigate how they perform when applied to full range editing, denoted as ”full” in Fig.6. It turns out that longer editing interval does not yield satisfying editing results. Especially for method like DiffusionCLIP, extending interval instead leads to loss of details and many artifacts. This observation proves again that improving the performance of editing based on diffusion model necessitates more than simply increasing the editing interval steps.

### 04 Jan 2024 v3

@@ Introduction
Figure 2: Overview of our proposed rectifier framework. The rectifier is a hypernetwork consisting of a global encoder and multiple subnet branches. It takes as input the original image $\bm{x}_{0}$ and the estimation at each step ($\mathbb{P}_{t}[\bm{\epsilon}_{t}^{\theta}(\bm{x}_{t})]$), targets to modulate the degraded residual features into offset weights, providing compensated information for high-fidelity reconstruction. We select the middle and up-sampling blocks of U-Net for modulate, considering that these blocks contain both high-level semantic information and low-level details. We also employ separable convolution to reduce the amount of generated parameters.
@@
While these methods point out how to compensate for the information gap, their reconstructions are still far from high-fidelity, and resorting to external network or classifier also makes it difficult to generalize to semantic editing tasks which mainly relies on diffusion’s internal representations. In this work, we propose a framework where a rectifier is incorporated to modulate residual features into offset weights, providing compensated information to help pretrained diffusion model achieving high-fidelity reconstructions. The framework is illustrated in Fig. 2. Our rectifier is a hypernetwork (David, Andrew, and Quoc 2016) which takes as input every step’s estimation and original image, expected to exploit the degradated residual features to fill the fidelity gap. The inputs first pass through a global encoder, then transformed by a series of sub-nets to generate layer-wise modulation. We choose to modulate the middle and up-sampling blocks of U-Net (Ronneberger, Fischer, and Brox 2015), considering they contain both high-level semantic information and low-level details. Furthermore, without the interference of other representations like classifier, our framework is highly adaptive for semantic editing tasks, and is easier to generalize to other diffusion-based downstream tasks.
@@
Another advantage of our editing training strategy is that we do not need to specify any heuristic-defined parameters to fit different attributes. It should be remarked here that for methods like Asyrp (Kwon, Jeong, and Uh 2022) and DiffusionCLIP (Kim, Kwon, and Ye 2022), neither of them employs editing through the entire denoising process. Asyrp halts editing prematurely and adds stochastic noise then to boost preceived quality, while DiffusionCLIP does not inverse images into complete noise for preserving their morphologies. Setting these parameters meticulously for every separate attribute is intricate and bothersome. Our method though, starts from pure noise and traverses the process throughly to get editing results, no extra settings are needed.
@@
In a comparable manner to reconstruction part, both quantitative and qualitative evaluation of editing are exhibited here too. Fig.5 presents some qualitative comparisons towards different methods trained under 50 inversion and sampling steps. Like previously noted, neither of Asyrp nor DiffusionCLIP employs edit through the entire denoising process, Asyrp applies stochastic noise during final process to boost perceived quality, and DiffusionCLIP begins editing from intermediate noisy images for preserving their morphologies. Our method, though starts from complete noise and traverses the whole denoising process to edit, still carries out editing results with exceptional quality. Several illustrative examples are, for instance, elements like hat and ring are kept intact during semantic editing, along with the preservation of image’s overall shape and background. Meanwhile, distortions or artifacts brought by conditional input text are avoided, and vital information loss is also alleviated.
@@
Doubts may arise regarding to whether our gain in editing comes from the rectifier or the training strategy. With the benefit of rectifier already proven in preceding part, we now focus on ablation studies to validate the effectiveness of our editing training strategy alone. Attribute ”smiling” is chosen for evaluating various algorithms. Fig.6 presents the performance of these algorithms. As shown, even without any enhancement from rectifier, our strategy still produces results with less distortion and information loss compared to others. This demonstrates our strategy could be employed as an independent and generalized training approach for editing tasks based on diffusion models. Furthermore, recall that neither Asyrp nor DiffusionCLIP implements editing through the entire process, they either stops prematurely or starts from incomplete noise. We thus investigate how they perform when applied to full range editing, denoted as ”full” in Fig.6. It turns out that longer editing interval does not yield satisfying editing results. Especially for method like DiffusionCLIP, extending interval instead leads to loss of details and many artifacts. This observation proves again that improving the performance of editing based on diffusion model necessitates more than simply increasing the editing interval steps.
@@ Acknowledgement
Supplementary Materials
A. Model Architectures
Our rectifier consists of a global encoder and a series of subnet branches to generate layer-wise modulations. ResNet34 (He et al. 2016) backbone is chosen as the global encoder. Inspired by how GAN-based methods cope with residual features (Alaluf et al. 2022; Pehlivan, Dalva, and Dundar 2023; Alaluf, Patashnik, and Cohen-Or 2021), we concatenate the original image and its estimation of every step $t$, constructing a 6-channel input to pass through global encoder. We also evaluate the performance of substracting the two inputs to obtain a 3-channel input, but the results are not good as 6-channel concatenation. The output is then processed by each subnet branch, which contains multiple convolutional layers to decrease the feature map size to 1 $\times$ 1, following with two convolutions to generate two slim matrices separately: $h\times w\times C_{in}\times 1$ and $h\times w\times 1\times C_{out}$ , the product of which is finally taken as the output modulated weight offset. Sinusoidal time embedding is integrated into each subnet branch using the same approach as U-Net, enabling the model with ability to perceive temporal representation. Detailed model architecture of rectifier and subnet are shown in Fig.1.
B. Training Details
During the training of editing, we set both $\lambda_{CLIP}$ and $\lambda_{recon}$ to 1. Adam optimizer is used for both reconstruction and editing training. We observe that the performance is substantially related to optimizer’s hyperparameter and learning schedule. We set weight decay to 1e-5, learning rate to 1e-3 and decreased by 0.9 every 5000 steps for reconstruction training. For editing, setting weight decay to 0, learning rate to 1e-3 and decreased by 0.9 every 10 steps works fine for most attributes. For those prompts that cause severe changes, we set learning schedule to lower decrease speed to avoid distortions. Only around 100 images are needed for training each attribute.
Figure 9: Detailed model architecture of rectifier (a) and its subnet (b).
C. Loss Function for Reconstruction Training
Like most of fintuning methods based on diffusion models, we also employ the original noise fitting loss used in diffusion model (Ho, Jain, and Abbeel 2020) itself for our reconstruction training. We also consider other loss functions like $\ell_{1}~{}loss$ which is commonly used in GAN’s high-fidelity fintuning tasks. Here we validate the effectiveness of different loss functions, including noise fitting loss $e~{}loss$, $\ell_{1}~{}loss$ and $\ell_{1}~{}loss$ with an output regularization $dw~{}loss$ (output delta weights should be as small as possible). Experiments are conducted under 50 inversion and sampling steps, the results of which are shown in Fig.2. From the figure we can see that $\ell_{1}~{}loss$ indeed restores the overall shape of input compared to original reconstruction, but the trade-off is that images become excessively smooth, along with loss of numerous details and texture, making the outcome not appeal like a real image (3rd row). The introduction of $dw~{}loss$ could alleviate this problem, as it confines the changes on original diffusion model, yet it also leads to failure of some expected restoration (4th row). Despite the other two, $e~{}loss$ achieves best results, for both recovering the overall shape and preserving the texture, it is surprisingly to see that $e~{}loss$ accomplish so much alone. Our exploration of loss function demonstrates that applying the loss function from previous methods (e.g. GAN-based methods) for finetuning diffusion models might not be the best choice, and the $e~{}loss$ used for training diffusion models may contain more information than expected.
Figure 10: Reconstruction under different loss functions. $e~{}loss$ accomplishes both restoring overall shapes and preserving details compared to other loss functions.
D. Hyper-parameters
Explorations are done towards the influence of different ranges of $\lambda_{CLIP}$ and $\lambda_{recon}$ in Eq. (6), as illustrated in Fig.3. The results demonstrate that larger $\lambda_{recon}$ tends to retain original content, while larger $\lambda_{CLIP}$ inclines to alter the attributes. Therefore, for attributes requiring large changes, larger $\lambda_{CLIP}$ is recommended, and larger $\lambda_{recon}$ works better for those which needs to maintain identity.
Figure 11: Exploration on hyper-parameters.
E. Editing Attributes Exploration
During editing, we indeed notice some attributes harder to train, resulting in subtle or irrelevant changes (e.g., angry). We reckon that this might due to the low frequency for such attributes occurred in the text used for CLIP training. To confirm this, we calculate the occurrence frequency of some attributes in LAION-400M (available): ’smile’: 198656, ’sad’: 199474, ’angry’: 21743. As a typical thorny attribute, angry is notably less frequent, providing some support to our hypothesis.
F. Additional Results
1. Reconstructions and editings comparisons on CelebA-HQ dataset (Karras et al. 2017) with representational based methods: Asyrp (Kwon, Jeong, and Uh 2022) and DiffusionCLIP (Kim, Kwon, and Ye 2022). Evaluations are conducted under various inversion and denoising steps from 50 to 1000, Fig.4, Fig.5 and Fig.6 represent these results separately. Both DiffusionCLIP and our method are trained under same 100 images, while official Asyrp checkpoints are loaded for evaluation.
2. Reconstructions and editing comparisons on AFHQ-Dogs (Choi et al. 2020) is shown in Fig.7. Besides two representational based editing methods mentioned above, we also test an image guidance based method GLIDE (Nichol et al. 2021) for AFHQ-Dogs dataset (pretrained GLIDE models on human faces are not released due to safety considerations), masks used in GLIDE inference are exhibited as well.
3. Image samples used in identity similarity calculation are displayed in Fig.8, we edit around 100 images with each method for evaluation.
4. More results on further applications including image-to-image translation and out-of-domain image editing are shown in Fig.9 and Fig.10, respectively.
Figure 12: Visual comparisons of reconstructions and editings, under 50 inversion and denoising steps.
Figure 13: Visual comparisons of reconstructions and editings, under 200 inversion and denoising steps. As the reconstruction quality for most images are quite well over 200 steps, this could also be seen as an ablation study toward the effectiveness of our editing training strategy alone.
Figure 14: Visual comparisons of reconstructions and editings under 1000 steps.
Figure 15: Visual comparisons of editings of AFHQ-Dogs. GLIDE achieves great detail preservation due to its mask mechanism, but sometimes causes inconsistency and unsatisfying results in semantic editing.
Figure 16: Image samples for calculating identity similarity.
Figure 17: Applications on image-to-image translation, compared to SDEdit(Meng et al. 2021). Every two columns of translation results share common random seed.
Figure 18: Applications on out-of-domain images editing on various attributes performed on METFACES (Karras et al. 2020).